When a ball is thrown upward at a speed of 19m/s, its height s above the ground (in meters) after t seconds is given by the formula
s=19t-24.01t
find the height of the ball after 3 seconds
I just can't figure out how to solve
Kim

Answer by Alan3354(69443)

When a ball is thrown upward at a speed of 19m/s, its height s above the ground (in meters) after t seconds is given by the formula
s=19t-24.01t
find the height of the ball after 3 seconds
---------
Sub 3 for t
s = 19*3 - 24.01*3 = -15.03 meters
Doesn't seem right, does it?
-----------------
Acutally, the equation is
s = 19t - 4.9t^2
at t=3
s = 19*3 - 4.9*9
s = 12.9 meters